OpenStudy (anonymous):

Solve the equation v + f – e = 2 for e.

OpenStudy (ikram002p):

ok then mean put e in a side alone ^_^

OpenStudy (anonymous):

yes ^

OpenStudy (ikram002p):

ok dude then solve huh !

OpenStudy (anonymous):

quack

OpenStudy (ikram002p):

ermmm just do this let e in side alone v + f – e = 2 e+2=v+f e=v+f-2

OpenStudy (anonymous):

i already got the answer man , thanks again though , little secret, i know all the answers to the questions i post , but i help others by giving them medals to the correct responses they give , such as the way you help someone out

OpenStudy (ikram002p):

lol haha :P